OpenStudy (kittiwitti1):

Chain Rule • "Find an equation of the tangent line at the given point." LaTeX below (2 questions).

OpenStudy (kittiwitti1):

\(\Large{\color{seagreen}{\text{A.}}}\) \(\large{y=3^{x^{2}}}\) at \((2,81)\) What I have so far:\[y'=(3^{x^{2}})'=(3^{u})'\times(x^{2})'=(3^{u}\ln{u})\times(2x)=(3^{x^{2}}\ln{x^{2}})\times(2x)\] \(\Large{\color{royalblue}{\text{B.}}}\) \(\large{y=\cos{\sin{x}}}\) at \((0,1)\) What I have so far:\[y'=(\cos{\sin{x}})'=(\cos{u})'\times(\sin{x})'=-\sin{(\sin{x})}\times\cos{x}\]

zepdrix (zepdrix):

Woops, \(\large\rm \left(3^x\right)'=3^x(ln3)\) Log of the base.

OpenStudy (kittiwitti1):

URGH! I made a typo. Sorry @zepdrix lol

OpenStudy (kittiwitti1):

Is it good otherwise, and is it necessary for me to proceed further?

zepdrix (zepdrix):

Well, yes, you've only finished about half of the problem so far :) Remember that the derivative function gives us `slope of our tangent line`. It doesn't give us the entire tangent line, only the slope.

OpenStudy (kittiwitti1):

Oh, I meant differential-wise haha. I want to make sure I'm solid enough to calculate the slope :-)

zepdrix (zepdrix):

Mmmm yes, looks good besides that! The way you wrote the function in part B is rather confusing. Use some brackets. cos(sin x). Derivative turned out well though.

OpenStudy (kittiwitti1):

Er, yes. I keep forgetting that in LaTeX lol

OpenStudy (kittiwitti1):

Alright, so I can start calculating the tangent slope in both?

zepdrix (zepdrix):

Sure.

OpenStudy (kittiwitti1):

Okay, well... For (B) I got this... when x=0,\[-\sin{(\sin{0})}\times\cos{0}=-\sin{0}\times1=0\]

OpenStudy (bonnieisflash1.0):

need help

OpenStudy (bonnieisflash1.0):

what do you need help on

OpenStudy (bonnieisflash1.0):

ok

OpenStudy (bonnieisflash1.0):

you can wait

OpenStudy (bonnieisflash1.0):

First solve for y as a function of x) and implicitly. 8. y2+2xy+4 = 0 at (-2, 2). In exercises 9-20, find the derivative y' (x) implicitly. In exercises 21-28, find an equation of the tangent line at the given point.

zepdrix (zepdrix):

For part B you got m=0? Ok that sounds right!

OpenStudy (kittiwitti1):

I have no idea what you are doing. @bonnieisflash1.0 Okay, thank you @zepdrix Except I am not sure how to find the equation of the tangent now. Would it be a horizontal line?

zepdrix (zepdrix):

If the slope is zero, then yes a horizontal line :)\[\large\rm y-y_o=m(x-x_o)\]Plugging in the point (0,1) and m=0,\[\large\rm y-1=0(x-0)\]

OpenStudy (kittiwitti1):

Ah, I see. Thank you for clarifying @zepdrix :-)

OpenStudy (kittiwitti1):

I got a very weird answer when putting the value \(x=2\) into the derived equation \((3^{x^{2}}\ln{3})\times2x\). It seems to be an overly large number... @zepdrix http://www.wolframalpha.com/input/?i=3%5E(4)ln(3)*4

OpenStudy (kittiwitti1):

\[\large{(3^{x^{2}}\ln{3})\times2x\rightarrow(3^{(2)^{2}}\ln{3})\times2(2)}\approx355.95...?\]

zepdrix (zepdrix):

It's difficult to see because the line is very very steep, https://www.desmos.com/calculator/zy17hydp3t But this slope value appears to be accurate, yes? I scaled the axes a bit so it's easier to see. 3^(x^2) is extreeeeeemely steep around x=2.

OpenStudy (kittiwitti1):

Hmm, yes, it does seem very steep... WAIT, how did you do that table on Desmos?! D:

zepdrix (zepdrix):

umm :d

zepdrix (zepdrix):

I dunno :d one of the buttons .. somewhere .. on there

OpenStudy (kittiwitti1):

D: ...so the slope is basically \(\approx356\)?

zepdrix (zepdrix):

Ya something like that, 4(ln3)3^4

OpenStudy (kittiwitti1):

Alright thank you :-) @zepdrix